A box contains 74 coins, only dimes and nickels. The amount of money in the box is $4.70. How many dimes and how many nickels are in the box?
Answer by addingup(3677)

d+n =74 subtract d from both sides:
n = 74-d
0.10d+0.05n = 4.70 substitute for n:
0.10d+0.05(74-d) = 4.70
0.10d+3.70-0.05d = 4.70
0.05d = 1
d = 20
n = 74-20 = 54